Three land lots have a TOTAL value of $96,000.
Lot A is $2,000 more than Lot B.
Lot B is $8,000 more than Lot C.
Q: What are the lot values for Lot A, B, and C?

Answer by info@email-tutors.com(9)

Hello:
Here is how we believe you should be able to solve the problem:
The total values of the lots is 96000. This means that a plus b plus c equals 96000: A + B + C = 96000
A = B+2000

Inserting this information into the equation, you get a different but equivalent equation:
(B + 2000) + B + C = 96000
However, B = C + 8000 transforming this information, we know that C = B - 8000
Inserting this information, you get:
(B + 2000) + B + (B - 8000)= 96000
This is equivalent to: 3B - 6000 = 96000
Add 6000 to both sides to get: 3B = 102000
Divide both sides by three to get: B = 34000
Using this value of B, you can find the rest of the numbers:
A = 34000 + 2000 = 36000
C = 34000 - 8000 = 26000
You can check your answer by adding: 36000 + 26000 + 34000
You will get 96000 as your answer.
We hope that we have been of help. If you need additional help with your homework, please check out our profile to visit our website: www.email-tutors.com. We offer email math tutoring at a rate of one dollar per problem.
Sincerely,
The Email-Tutors dot com Team